OpenStudy (anonymous):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.) 2 sin^2 θ − sin θ − 1 = 0

OpenStudy (turingtest):

we can factor this: \[(2\sin \theta+1)(\sin \theta-1)=0\]\[\sin \theta={-1\over2}\to \theta=\left\{ {-\pi \over6}+2k \pi,{7\pi \over6}+2k \pi \right\}\]\[\sin \theta=1 \to \theta=\left\{ {\pi \over2}+2k \pi \right\}\]